Carl and Dennie's are brothers. The sum of their ages is 29. In 6 year's time Carl will be same age as Dennis was 3 years ago. Write down the appropriate equation.
Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39618)
Answer by greenestamps(13200)



Here is ONE way to set up the problem for solving....

let x = Carl's age
then 29-x = Dennie's age

Carl's age 6 years from now will be x+6
Dennie's age 3 years ago was (29-x)-3 = 26-x

In 6 year's time Carl will be same age as Dennie was 3 years ago:

x+6 = 26-x

That is ONE equation that can be used to solve the problem.

Setting the problem up differently might lead to a very different equation....